# Nonasymptotic Analysis of Accelerated Methods With Inexact Oracle Under Absolute Error Bound

### v1

@@ 3.2 PEP-inspired convergence analysis of iGOGM
$P_{k,k}\triangleq$ $\frac{(2\alpha_{k+1}\alpha_{k}+A_{k})^{2}}{4L(A_{k+1}-\alpha_{k+1}^{2})}+\sum_{j=k+1}^{K-1}\frac{\alpha_{k}^{2}\alpha_{j+1}^{2}}{L(A_{j+1}-\alpha_{j+1}^{2})},$
$P_{k,i}\triangleq$ $\frac{\alpha_{k+1}\alpha_{i}(A_{k}+2\alpha_{k+1}\alpha_{k})}{L(A_{k+1}-\alpha_{k+1}^{2})}+\sum_{j=k+1}^{K-1}\frac{2\alpha_{j+1}^{2}\alpha_{k}\alpha_{i}}{L(A_{j+1}-\alpha_{j+1}^{2})}.$
Hence, the above equation can be rewritten as
$\sum_{k=0}^{K-1}P_{k,k}\left\|\mathbf{e}_{k}\right\|^{2}+\sum_{k=0}^{K-1}\sum_{i=0}^{k-1}P_{k,i}\left\langle\mathbf{e}_{k},\mathbf{e}_{i}\right\rangle$
$=$ $-\sum_{k=0}^{K-1}\sum_{i=0}^{k-1}\frac{1}{2}P_{k,i}\left\|\mathbf{e}_{k}-\mathbf{e}_{i}\right\|^{2}+\sum_{k=0}^{K-1}\sum_{i=0}^{k-1}\frac{1}{2}P_{k,i}(\left\|\mathbf{e}_{k}\right\|^{2}+\left\|\mathbf{e}_{i}\right\|^{2})+\sum_{k=0}^{K-1}P_{k,k}\left\|\mathbf{e}_{k}\right\|^{2}$
$=$ $-\sum_{k=0}^{K-1}\sum_{i=0}^{k-1}\frac{1}{2}P_{k,i}\left\|\mathbf{e}_{k}-\mathbf{e}_{i}\right\|^{2}+\sum_{k=0}^{K-1}\sum_{i=0}^{k-1}\frac{1}{2}P_{k,i}\left\|\mathbf{e}_{k}\right\|^{2}+\sum_{k=0}^{K-1}\sum_{i=k+1}^{K-1}\frac{1}{2}P_{i,k}\left\|\mathbf{e}_{k}\right\|^{2}+\sum_{k=0}^{K-1}P_{k,k}\left\|\mathbf{e}_{k}\right\|^{2}$
$\stackrel{{\scriptstyle(iii)}}{{\leq}}$ $\sum_{k=0}^{K-1}\left(P_{k,k}+\frac{1}{2}\sum_{i=0}^{k-1}P_{k,i}+\frac{1}{2}\sum_{i=k+1}^{K-1}P_{i,k}\right)\left\|\mathbf{e}_{k}\right\|^{2}.$
@@
$=$ $\frac{A_{k}(1+2\alpha_{k+1})(A_{k}+2\alpha_{k}\alpha_{k+1})}{4L(A_{k+1}-\alpha_{k+1}^{2})}+\sum_{i=k+1}^{K-1}\frac{A_{i}(1+2\alpha_{i+1})\alpha_{k}\alpha_{i+1}}{2L(A_{i+1}-\alpha_{i+1}^{2})}.$
Inserting these results into the original inequality, we get the final bound. ∎
The key steps of the proof are the equalities $(i)$, $(ii)$, and inequality $(iii)$. The equality $(i)$ can be seen as the step of $\mathbf{M}=\frac{1}{\tau}\mathbf{w}_{\tau}\mathbf{w}_{\tau}^{\top}+\mathbf{R}$ if one uses the Gram matrix $\mathbf{G}$ and standard basis vectors to rewrite this equality. The equality $(ii)$ can be seen as the reformulation of the Shur complement with the block $[\frac{A_{k+1}-\alpha_{k+1}^{2}}{LA_{K}}]$ of $\mathbf{R}^{\backslash[1,K+2]}$. The inequality (iii) is indeed the step of finding $\{u_{k}\}$ such that the matrix $\mathbf{S}$ is diagonally dominant. This observation shows that the steps we use to find and verify the feasible solution ($\mathbf{M}\succeq 0$) can be used to generate the theoretical proof.
When $\left\|\mathbf{e}_{k}\right\|\equiv 0$, the error term will be zero and the upper bound recovers the result of the exact GOGM. When $\left\|\mathbf{e}_{k}\right\|\neq 0$, this term is only determined by the total iteration number $K$ and the stepsize $\{\alpha_{k}\}$. It is independent of the initial condition $\left\|\mathbf{x}_{0}-\mathbf{x}_{*}\right\|$ and can be calculated when $\{\alpha_{k}\}$ is fixed.
@@
The inexactness assumption and corresponding analysis in [49] can also handle unbounded feasible regions. However, their inexactness assumption for convex problems requires the gradient estimate to be a subgradient, which is not a requirement in our work. Furthermore, their inexactness level along iterations has to be the same. Finally, in their FGM analysis, the accumulated error depends on the initial condition $R$. In contrast, our approach provides a pure separation of the convergence rate with the exact gradient oracle and the accumulated error. Our error term is independent of $R$ and holds for any arbitrary error along iterations.
3.3 PEP-inspired convergence analysis of iGFGM
Theorem 3.5.
The sequence generated by Algorithm 6 satisfies
4 Exploiting the rate-error tradeoff and finding the optimal inexactness schedule
In Theorem 3.3, we provided the convergence bound of iGOGM with the explicit expression for the accumulated error. This section first looks into the tradeoff between the convergence rate and accumulated error through the stepsize with a fixed gradient inexactness level along iterations. Second, we aim to optimize the inexactness levels across iterations, for a given stepsize, so that the total effort to obtain the gradient oracles is minimized.
4.1 Convergence rate and the accumulated error tradeoff
Considering the convergence bound $\frac{L\left\|\mathbf{x}_{0}-\mathbf{x}_{*}\right\|^{2}}{4A_{K}}+\sum_{k=0}^{K-1}u_{k}b_{k}^{2}$, with $\{A_{k}\},\{u_{k}\}$ being functions of $\{\alpha_{k}\}$, one can propose to minimize the bound with respect to $\{\alpha_{k}\}$. However, we note that such a problem is a high-order polynomial optimization and its solution is intractable in general.
Even if one is satisfied with the numerical solution to find the optimized algorithm, then it is preferable to solve the minimax problem (9) following the relaxation approach in [22] which we also used in Appendix B. This procedure generates the optimized stepsizes and the resulting convergence bound is better than the solution of the polynomial optimization problem as the derived theoretical $\{u_{i}\}$ in Lemma 3.2 is not necessarily the optimal solution.
Here we narrow down the iGOGM to iOGM-a, with the stepsize $\alpha_{k}=\frac{i+a}{a}$ and $A_{k}=\sum_{i=0}^{k}\frac{i+a}{a}=\frac{(k+2a)(k+1)}{2a}$. The condition $A_{k}>\alpha_{k}^{2}$ holds for any $k$ when $a>2$. Note that when $K\gg a$, the convergence rate is still $\mathcal{O}(K^{-2})$ in the exact case. Furthermore, [42] shows that the exact algorithm OGM-a has the asymptotic worst-case bound for function value as $\frac{a}{2}K^{-2}$ and the smallest gradient norm as $\frac{a\sqrt{6}}{2\sqrt{a-2}}K^{-1.5}$; it achieves the best performance of these two measures simultaneously when $a=4$. Both the simplicity of the stepsize and the good performances over function value and gradient norm indicate that OGM-a is a good selection under exact oracles and is worth further study for the inexact oracles too.
Replacing the stepsizes of OGM-$a$ in Theorem 3.3, we get the coefficients of the error as
$+\sum_{i=k+1}^{K-1}\frac{(k+a)(2i^{3}+(9a+4)i^{2}+(13a^{2}+13a+2)i+6a^{3}+10a^{2}+4a}{La(K+1)(K+2a)((a-2)i+2a^{2}-2a-2)},$
which can be further simplified when $a$ is assigned with a value.
Take $a=4$ as an example, which has both best convergence rates for function value and the gradient norm under exact oracle, the $u_{k}$ is
Assume $b_{k}\equiv b$, then the accumulated error is calculated as
$b^{2}\sum_{k=0}^{K-1}u_{k}=\frac{b^{2}K(12K^{3}+303K^{2}+2687K+8758)}{480L(K+8)}.$
The above results indicate that the accumulated error is in $\mathcal{O}(\frac{b^{2}K^{3}}{L})$.
For any arbitrary $a$ values, it is difficult to derive a closed form of $u_{k}$ since the denominator can only be simplified as $k+c$ structure when $a$ is given as a specific number. Hence, we provide a numerical result of $u_{k}$ for given $K$ with different $a$ in Figure 4. For the same $a$ and $k$, bigger $K$ results in bigger $u_{k}$. For all $a$ values, $\{u_{k}\}$ first increases and then decreases. For a fixed $k$ and $K$, bigger $a$ results in smaller $u_{k}$ values.
Figure 4: Values of $u_{k}$ for OGM-$a$ with different values of $a$ and $K$.
Based on this result, to reduce the accumulated error, one can consider increasing $a$. However, it will adversely slow down the convergence rate as $A_{K}=\frac{(K+2a)(K+1)}{2a}$ and decreases when $a$ is increased. Under this circumstance, we provide the comparison of convergence rate and accumulated error with different iteration numbers $K$, and $a$ in Figure 5, where $\tau=\frac{L}{4A_{K}}$. The convergence rate $\tau$ is smaller when $a$ is smaller while the accumulated error is bigger. When $a=4$, $\sum_{k=0}^{K-1}u_{k}$ is between $K^{2}$ and $K^{3}$, better than our relaxed analysis result. Another interesting fact is that when $K\ll a$ ($a=10^{5}$ and $a=10^{6}$), the convergence rate $\tau=\mathcal{O}(K^{-1})$ and the accumulate error is also $\mathcal{O}(K^{1})$. If $K\ll a$, we have $\alpha_{k}=1$, and the proposition below provides the convergence bound of this extreme scenario.
Proposition 4.1.
4.2 Optimal inexactness schedule
Since the derived $u_{k}$ is not constant, it will be beneficial to set $b_{k}$ based on the value of $u_{k}$. Recall the definition of $\mathbb{\eta}$ from (2). When $r=1$, we can aim to optimize over the total $\eta$-complexity, i.e., $\sum_{k=0}^{K-1}\eta_{k}$, so that the accumulated error will not deteriorate the convergence rate. The problem is given as
[…]
$\operatorname{s.t.}\quad$ $\sum_{k=0}^{K-1}u_{k}b_{k}^{2}\leq\frac{LR^{2}}{4A_{K}}.$
This problem is a convex optimization when $h$ is convex and decreasing over linear constraint and at its optimal solution, the constraint is binding. Hence, the inequality constraint can be replaced with the equality one. Let $\lambda$ be the dual multiplier of the constraint, the Lagrangian function is
[…]
Lemma 4.2 ($h(\eta)$ with power law decay).
Furthermore, the optimal primal and dual solutions of (OPT-b) are
$b_{k}^{*}=\frac{\sqrt{L}R}{2\sqrt{A_{K}\sum_{k=0}^{K-1}u_{k}^{1/(1+2c_{2})}}u_{k}^{d/(1+2c_{2})}},\quad\lambda^{*}=\frac{c^{1/c_{2}}}{2c_{2}}\left(\frac{LR^{2}}{4A_{K}}\right)^{-(1+2c_{2})/(2c_{2})}\left(\sum_{k=0}^{K-1}u_{k}^{1/(1+2c_{2})}\right)^{(1+2c_{2})/(2c_{2})}.$
Proof.
With power law decay, (13) is written as
$-\frac{c^{1/c_{2}}(b_{k}^{*})^{-(1+c_{2})/c_{2}}}{c_{2}}+2\lambda^{*}b_{k}^{*}u_{k}=0,\quad k=0,1,\cdots,K-1.$
From this equation, $b_{k}^{*}$ can be calculated as
[…]
Inserting this equation into (13), and solve for $\lambda^{*}$ we get
Replacing this solution into (15), the final result can be derived. ∎
Lemma 4.3 ($h(\eta)$ with exponential decay).
Let $h(\eta)=q_{1}q_{2}^{-\eta}$ with $q_{1}>0$ and $q_{2}>1$ be two constants. Then
and, the optimal primal and dual solutions are
Proof.
Similar to the proof of Lemma 4.2, from (13) we have
$u_{k}(b_{k}^{*})^{2}=\frac{1}{2\lambda^{*}\log q_{2}}.$ (16)
Inserting this into (14),
$\lambda^{*}=\frac{2KA_{K}}{LR^{2}\log q_{2}}.$
By the previous two equations, we can derive the final solution. ∎
To quantify the improvement after the optimized inexactness level, we calculate the total $\eta$-complexity for iOGM-4 algorithm for the two structures of $h(\eta)$ functions. We fix $L=1$ and $R=1$, then solve the constant $b$ through
$\bar{b}^{2}\sum_{k=0}^{K-1}\leq\frac{LR^{2}}{4A_{K}},$
and calculate the optimized $b$ by the solutions we derived in Lemmas 4.2 and 4.3. In Figure 6, we report the result for power law decay with four different values of $c_{2}$. Note that the $\eta$-complexity decreases with increased $c_{2}$. If we focus on the improvement of $\eta$-complexity for the optimized $\{b_{k}\}$, the improvement is in the same order as the complexity of constant $b$, which shows optimizing over $b$ can decrease the $\eta$-complexity significantly. As for the exponential decay in Figure 7, the improvement is not as large as that of the power law decay. Although the improvement is not as large as that of the power law decay, considering its absolute value, it still represents a considerable improvement.
@@ 5 Conclusions
In this paper, we analyze the Generalized Optimal Gradient Method (GOGM) with the inexact gradient oracle under the absolute error assumption. By utilizing the Performance Estimation Problem (PEP) analysis tool, a new upper bound for the convergence rate of GOGM is derived. This bound is composed of two independent components, the convergence rate obtained from the exact gradient oracle and the accumulated error resulting from the gradient inexactness. Such a bound demonstrates the effect of the inexactness and does not require boundedness of the feasible region. Furthermore, this bound is independent of the initial condition.
From this convergence bound, the optimal schedule to set oracle inexactness along iterations is proposed. Such a study aims to minimize the sampling or computational effort for gradient estimation while maintaining the convergence rate.
@@ Appendix A Supplementary lemmas and proofs
A.1 Proof of composition optimization
@@
Assume for any $\mathbf{v}$ and $\mathbf{x}$, $\left\|\nabla f(\mathbf{v})\right\|\leq C_{f}$ and $\left\|\nabla\mathbf{g}(\mathbf{x})\right\|\leq C_{g}$. Also, $f(\mathbf{v})$ is Lipschitz smooth with constant $L_{f}$, $\mathbb{E}_{i}[\left\|\nabla\mathbf{g}_{i}(\mathbf{x})-\nabla\mathbf{g}(\mathbf{x})\right\|^{2}]\leq\sigma^{2}_{\nabla}$ and $\mathbb{E}[\left\|\mathbf{g}_{i}(\mathbf{x})-\mathbf{g}(\mathbf{x})\right\|^{2}]\leq\sigma^{2}_{\mathbf{g}}$. Assume $\left|\mathcal{S}_{\nabla}\right|$ and $\left|\mathcal{S}_{\mathbf{g}}\right|$ are properly chosen so $\frac{2C_{f}^{2}}{b^{2}\left|\mathcal{S}_{\nabla}\right|}\sigma_{\nabla}^{2}+\frac{2C_{g}^{2}L_{f}^{2}}{b^{2}\left|\mathcal{S}_{\mathbf{g}}\right|}\sigma_{\mathbf{g}}^{2}\leq 1$, then the error of the estimated gradient is bounded as
~$\left\|\tilde{\nabla}\mathbf{g}(\mathbf{x})\nabla f(\tilde{\mathbf{g}}(\mathbf{x}))-\nabla\mathbf{g}(\mathbf{x})\nabla f(\mathbf{g}(\mathbf{x}))\right\|^{2}\leq b^{2},$
with probability at least $1-\frac{2C_{f}^{2}}{b^{2}\left|\mathcal{S}_{\nabla}\right|}\sigma_{\nabla}^{2}+\frac{2C_{g}^{2}L_{f}^{2}}{b^{2}\left|\mathcal{S}_{\mathbf{g}}\right|}\sigma_{\mathbf{g}}^{2}$.
@@
~$\left\|\tilde{\nabla}\mathbf{g}(\mathbf{x})\nabla f(\tilde{\mathbf{g}}(\mathbf{x}))-\nabla\mathbf{g}(\mathbf{x})\nabla f(\mathbf{g}(\mathbf{x}))\right\|^{2}$
$=$ ~$\left\|\tilde{\nabla}\mathbf{g}(\mathbf{x})\nabla f(\tilde{\mathbf{g}}(\mathbf{x}))-\nabla\mathbf{g}(\mathbf{x})\nabla f(\tilde{\mathbf{g}}(\mathbf{x}))+\nabla\mathbf{g}(\mathbf{x})\nabla f(\tilde{\mathbf{g}}(\mathbf{x}))-\nabla\mathbf{g}(\mathbf{x})\nabla f(\mathbf{g}(\mathbf{x}))\right\|^{2}$
$\leq$ $2\left\|\tilde{\nabla}\mathbf{g}(\mathbf{x})\nabla f(\tilde{\mathbf{g}}(\mathbf{x}))-\nabla\mathbf{g}(\mathbf{x})\nabla f(\tilde{\mathbf{g}}(\mathbf{x}))\right\|^{2}+2\left\|\nabla\mathbf{g}(\mathbf{x})\nabla f(\tilde{\mathbf{g}}(\mathbf{x}))-\nabla\mathbf{g}(\mathbf{x})\nabla f(\mathbf{g}(\mathbf{x}))\right\|^{2}$
$\leq$ $2\left\|\nabla f(\tilde{\mathbf{g}}(\mathbf{x}))\right\|^{2}\left\|\tilde{\nabla}\mathbf{g}(\mathbf{x})-\nabla\mathbf{g}(\mathbf{x})\right\|^{2}+2\left\|\nabla\mathbf{g}(\mathbf{x})\right\|^{2}\left\|\nabla f(\tilde{\mathbf{g}}(\mathbf{x}))-\nabla f(\mathbf{g}(\mathbf{x}))\right\|^{2}$
$\leq$ $2C_{f}^{2}\left\|\tilde{\nabla}\mathbf{g}(\mathbf{x})-\nabla\mathbf{g}(\mathbf{x})\right\|^{2}+2C_{g}^{2}L_{f}^{2}\left\|\tilde{\mathbf{g}}(\mathbf{x})-\mathbf{g}(\mathbf{x})\right\|^{2}.$
@@
$\mathbb{E}[\left\|\tilde{\nabla}\mathbf{g}(\mathbf{x})\nabla f(\tilde{\mathbf{g}}(\mathbf{x}))-\nabla\mathbf{g}(\mathbf{x})\nabla f(\mathbf{g}(\mathbf{x}))\right\|^{2}]$
$\leq$ $\frac{2C_{f}^{2}}{\left|\mathcal{S}_{\nabla}\right|}\sigma_{\nabla}^{2}+\frac{2C_{g}^{2}L_{f}^{2}}{\left|\mathcal{S}_{\mathbf{g}}\right|}\sigma_{\mathbf{g}}^{2}.$

### v2

@@ 4.1 PEP-inspired proof of the convergence analysis of iGOGM
$P_{k,k}\triangleq$ $\frac{(2\alpha_{k+1}\alpha_{k}+A_{k})^{2}}{4LA_{K}(A_{k+1}-\alpha_{k+1}^{2})}+\sum_{j=k+1}^{K-1}\frac{\alpha_{k}^{2}\alpha_{j+1}^{2}}{LA_{K}(A_{j+1}-\alpha_{j+1}^{2})},$
$P_{k,i}\triangleq$ $\frac{\alpha_{k+1}\alpha_{i}(A_{k}+2\alpha_{k+1}\alpha_{k})}{LA_{K}(A_{k+1}-\alpha_{k+1}^{2})}+\sum_{j=k+1}^{K-1}\frac{2\alpha_{j+1}^{2}\alpha_{k}\alpha_{i}}{LA_{K}(A_{j+1}-\alpha_{j+1}^{2})}.$
Hence, the above inequality can be rewritten as
$f(\mathbf{x}_{K})-f_{*}-\frac{1}{2L}\left\|\nabla f(\mathbf{x}_{K})\right\|^{2}$
$\leq$ $\frac{LR^{2}}{4A_{K}}+\sum_{k=0}^{K-1}u_{k}(b_{k}^{2}-\left\|\mathbf{e}_{k}\right\|^{2})+\sum_{k=0}^{K-1}P_{k,k}\left\|\mathbf{e}_{k}\right\|^{2}+\sum_{k=0}^{K-1}\sum_{i=0}^{k-1}P_{k,i}\left\langle\mathbf{e}_{k},\mathbf{e}_{i}\right\rangle$
$=$ $\frac{LR^{2}}{4A_{K}}+\sum_{k=0}^{K-1}u_{k}(b_{k}^{2}-\left\|\mathbf{e}_{k}\right\|^{2})-\sum_{k=0}^{K-1}\sum_{i=0}^{k-1}\frac{1}{2}P_{k,i}\left\|\mathbf{e}_{k}-\mathbf{e}_{i}\right\|^{2}$
$+\sum_{k=0}^{K-1}\sum_{i=0}^{k-1}\frac{1}{2}P_{k,i}(\left\|\mathbf{e}_{k}\right\|^{2}+\left\|\mathbf{e}_{i}\right\|^{2})+\sum_{k=0}^{K-1}P_{k,k}\left\|\mathbf{e}_{k}\right\|^{2}$
$=$ $\frac{LR^{2}}{4A_{K}}+\sum_{k=0}^{K-1}u_{k}(b_{k}^{2}-\left\|\mathbf{e}_{k}\right\|^{2})-\sum_{k=0}^{K-1}\sum_{i=0}^{k-1}\frac{1}{2}P_{k,i}\left\|\mathbf{e}_{k}-\mathbf{e}_{i}\right\|^{2}$
$+\sum_{k=0}^{K-1}\sum_{i=0}^{k-1}\frac{1}{2}P_{k,i}\left\|\mathbf{e}_{k}\right\|^{2}+\sum_{k=0}^{K-1}\sum_{i=k+1}^{K-1}\frac{1}{2}P_{i,k}\left\|\mathbf{e}_{k}\right\|^{2}+\sum_{k=0}^{K-1}P_{k,k}\left\|\mathbf{e}_{k}\right\|^{2}$
$\leq$ $\frac{LR^{2}}{4A_{K}}+\sum_{k=0}^{K-1}u_{k}b_{k}^{2}+\sum_{k=0}^{K-1}\left(P_{k,k}+\frac{1}{2}\sum_{i=0}^{k-1}P_{k,i}+\frac{1}{2}\sum_{i=k+1}^{K-1}P_{i,k}-u_{k}\right)\left\|\mathbf{e}_{k}\right\|^{2}.$
@@
$=$ $\frac{A_{k}(1+2\alpha_{k+1})(A_{k}+2\alpha_{k}\alpha_{k+1})}{4LA_{K}(A_{k+1}-\alpha_{k+1}^{2})}+\sum_{i=k+1}^{K-1}\frac{A_{i}(1+2\alpha_{i+1})\alpha_{k}\alpha_{i+1}}{2LA_{K}(A_{i+1}-\alpha_{i+1}^{2})}.$
If each $u_{k}$ is large enough, in this case,
$u_{k}\geq\frac{A_{k}(1+2\alpha_{k+1})(A_{k}+2\alpha_{k}\alpha_{k+1})}{4LA_{K}(A_{k+1}-\alpha_{k+1}^{2})}+\sum_{i=k+1}^{K-1}\frac{A_{i}(1+2\alpha_{i+1})\alpha_{k}\alpha_{i+1}}{2LA_{K}(A_{i+1}-\alpha_{i+1}^{2})},$
then the coefficients of $\left\|\mathbf{e}_{k}\right\|^{2}$ will be negative and we can derive the convergence result of iGOGM as
$f(\mathbf{x}_{K})-f_{*}-\frac{1}{2L}\left\|\nabla f(\mathbf{x}_{K})\right\|^{2}\leq\frac{LR^{2}}{4A_{K}}+\sum_{k=0}^{K-1}u_{k}b_{k}^{2}.$
The selection of the lower bound of $u_{k}$ is equivalent to establishing conditions for $\mathbf{S}$ in (18) to be a diagonally dominated matrix. Inserting these results into the original inequality, we get the final bound. ∎
@@
Note that Theorem 2.2 is valid only for $0<\alpha_{k}^{2}<A_{k}$, i.e., $\alpha_{k}^{2}$ cannot be equal to $A_{k}$, which reflects a limitation of the analytically feasible solution we derived. However, as demonstrated in Theorem 2.4, for iFGM, we can set $\alpha_{k}^{2}=A_{k}$. This particular choice allows iFGM to attain its fastest convergence rate if the oracle is exact.
@@ 5 Conclusions
In this paper, we analyze the Generalized Optimized Gradient Method (GOGM) with the inexact gradient oracle, i.e., iGOGM, under the absolute error assumption. By utilizing the Performance Estimation Problem (PEP) analysis tool, a new upper bound for the convergence rate of the iGOGM algorithm is derived. This bound comprises two components: the convergence rate obtained from the exact gradient oracle and the accumulated error resulting from the gradient oracle’s inexactness. Such a bound demonstrates the effect of the inexactness and does not require boundedness of the feasible region. Furthermore, the accumulated error component of the bound is independent of the initial condition.
Furthermore, from this convergence bound, the optimal schedule to set the oracle inexactness along iterations is proposed. Such a study aims to minimize the sampling or computational effort for gradient estimation while maintaining the convergence rate.
@@ Appendix A Supplementary lemmas and proofs
A.1 Details on motivational examples
Bilevel optimization.
Consider the bilevel optimization problem
$\min_{\mathbf{x}}$ $\quad f(\mathbf{x};\mathbf{y}^{*}(\mathbf{x}))$
$\left\|\tilde{\nabla}f(\mathbf{x};\tilde{\mathbf{y}}(\mathbf{x}))-\nabla f(\mathbf{x};\mathbf{y}^{*}(\mathbf{x}))\right\|\leq C\left\|\tilde{\mathbf{y}}(\mathbf{x})-\mathbf{y}^{*}(\mathbf{x})\right\|,$
Now let $\tilde{\mathbf{y}}_{k}(\mathbf{x})$ be the solution of the lower-level problem by the gradient descent method with stepsize $\frac{2}{\mu_{g}+L_{g}}$ after $k$ iterations. From the iteration complexity of GD for the class $\mathcal{F}_{\mu,L}$, the gradient inexactness can be bounded as
$\left\|\tilde{\nabla}f(\mathbf{x};\tilde{\mathbf{y}}_{k}(\mathbf{x}))-\nabla f(\mathbf{x};\mathbf{y}^{*}(\mathbf{x}))\right\|\leq C\left(\frac{Q_{g}-1}{Q_{g}+1}\right)^{k}\left\|\mathbf{y}_{0}-\mathbf{y}^{*}(\mathbf{x})\right\|,$
Composition optimization.
Consider the nested composition optimization
Assume for any $\mathbf{v}$ and $\mathbf{x}$, $\left\|\nabla h(\mathbf{v})\right\|\leq C_{h}$ and $\left\|\nabla\mathbf{g}(\mathbf{x})\right\|\leq C_{g}$. Furthermore, assume $h(\mathbf{v})$ is Lipschitz smooth with constant $L_{h}$, $\mathbb{E}_{i}[\left\|\nabla\mathbf{g}_{i}(\mathbf{x})-\nabla\mathbf{g}(\mathbf{x})\right\|^{2}]\leq\sigma^{2}_{\nabla}$, and $\mathbb{E}[\left\|\mathbf{g}_{i}(\mathbf{x})-\mathbf{g}(\mathbf{x})\right\|^{2}]\leq\sigma^{2}_{\mathbf{g}}$. Then with probability at least $1-\frac{1}{\epsilon}$, the inexact gradient estimate is bounded as
$\left\|\tilde{\nabla}\mathbf{g}(\mathbf{x})\nabla h(\tilde{\mathbf{g}}(\mathbf{x}))-\nabla\mathbf{g}(\mathbf{x})\nabla h(\mathbf{g}(\mathbf{x}))\right\|^{2}\leq\epsilon\left(\frac{2C_{h}^{2}}{\left|\mathcal{S}_{\nabla}\right|}\sigma_{\nabla}^{2}+\frac{2C_{g}^{2}L_{h}^{2}}{\left|\mathcal{S}_{\mathbf{g}}\right|}\sigma_{\mathbf{g}}^{2}\right).$
@@ A.1 Details on motivational examples
Assume for any $\mathbf{v}$ and $\mathbf{x}$, $\left\|\nabla h(\mathbf{v})\right\|\leq C_{h}$ and $\left\|\nabla\mathbf{g}(\mathbf{x})\right\|\leq C_{g}$. Also, $f(\mathbf{v})$ is Lipschitz smooth with constant $L_{h}$, $\mathbb{E}_{i}[\left\|\nabla\mathbf{g}_{i}(\mathbf{x})-\nabla\mathbf{g}(\mathbf{x})\right\|^{2}]\leq\sigma^{2}_{\nabla}$ and $\mathbb{E}[\left\|\mathbf{g}_{i}(\mathbf{x})-\mathbf{g}(\mathbf{x})\right\|^{2}]\leq\sigma^{2}_{\mathbf{g}}$. Assume $\left|\mathcal{S}_{\nabla}\right|$ and $\left|\mathcal{S}_{\mathbf{g}}\right|$ are properly chosen so $\frac{2C_{h}^{2}}{b^{2}\left|\mathcal{S}_{\nabla}\right|}\sigma_{\nabla}^{2}+\frac{2C_{g}^{2}L_{h}^{2}}{b^{2}\left|\mathcal{S}_{\mathbf{g}}\right|}\sigma_{\mathbf{g}}^{2}\leq 1$, then the error of the estimated gradient is bounded as
$\left\|\tilde{\nabla}\mathbf{g}(\mathbf{x})\nabla h(\tilde{\mathbf{g}}(\mathbf{x}))-\nabla\mathbf{g}(\mathbf{x})\nabla h(\mathbf{g}(\mathbf{x}))\right\|^{2}\leq b^{2},$
with probability at least $1-\frac{2C_{h}^{2}}{b^{2}\left|\mathcal{S}_{\nabla}\right|}\sigma_{\nabla}^{2}+\frac{2C_{g}^{2}L_{h}^{2}}{b^{2}\left|\mathcal{S}_{\mathbf{g}}\right|}\sigma_{\mathbf{g}}^{2}$.
@@
$\left\|\tilde{\nabla}\mathbf{g}(\mathbf{x})\nabla h(\tilde{\mathbf{g}}(\mathbf{x}))-\nabla\mathbf{g}(\mathbf{x})\nabla h(\mathbf{g}(\mathbf{x}))\right\|^{2}$
$=$ $\left\|\tilde{\nabla}\mathbf{g}(\mathbf{x})\nabla h(\tilde{\mathbf{g}}(\mathbf{x}))-\nabla\mathbf{g}(\mathbf{x})\nabla h(\tilde{\mathbf{g}}(\mathbf{x}))+\nabla\mathbf{g}(\mathbf{x})\nabla h(\tilde{\mathbf{g}}(\mathbf{x}))-\nabla\mathbf{g}(\mathbf{x})\nabla h(\mathbf{g}(\mathbf{x}))\right\|^{2}$
$\leq$ $2\left\|\tilde{\nabla}\mathbf{g}(\mathbf{x})\nabla h(\tilde{\mathbf{g}}(\mathbf{x}))-\nabla\mathbf{g}(\mathbf{x})\nabla h(\tilde{\mathbf{g}}(\mathbf{x}))\right\|^{2}+2\left\|\nabla\mathbf{g}(\mathbf{x})\nabla h(\tilde{\mathbf{g}}(\mathbf{x}))-\nabla\mathbf{g}(\mathbf{x})\nabla h(\mathbf{g}(\mathbf{x}))\right\|^{2}$
$\leq$ $2\left\|\nabla h(\tilde{\mathbf{g}}(\mathbf{x}))\right\|^{2}\left\|\tilde{\nabla}\mathbf{g}(\mathbf{x})-\nabla\mathbf{g}(\mathbf{x})\right\|^{2}+2\left\|\nabla\mathbf{g}(\mathbf{x})\right\|^{2}\left\|\nabla h(\tilde{\mathbf{g}}(\mathbf{x}))-\nabla h(\mathbf{g}(\mathbf{x}))\right\|^{2}$
$\leq$ $2C_{h}^{2}\left\|\tilde{\nabla}\mathbf{g}(\mathbf{x})-\nabla\mathbf{g}(\mathbf{x})\right\|^{2}+2C_{g}^{2}L_{h}^{2}\left\|\tilde{\mathbf{g}}(\mathbf{x})-\mathbf{g}(\mathbf{x})\right\|^{2}.$
@@
$\mathbb{E}[\left\|\tilde{\nabla}\mathbf{g}(\mathbf{x})\nabla h(\tilde{\mathbf{g}}(\mathbf{x}))-\nabla\mathbf{g}(\mathbf{x})\nabla h(\mathbf{g}(\mathbf{x}))\right\|^{2}]$
$\leq$ $\frac{2C_{h}^{2}}{\left|\mathcal{S}_{\nabla}\right|}\sigma_{\nabla}^{2}+\frac{2C_{g}^{2}L_{h}^{2}}{\left|\mathcal{S}_{\mathbf{g}}\right|}\sigma_{\mathbf{g}}^{2}.$